The current in a 70.0-mH inductor changes with time as i = 2.00ẹ-7.00t, where i is in amperes and t is in seconds. (a) Find the magnitude of the induced emf at t 1.00 s mV (b) Find the magnitude of the induced emf at t- 4.00 s. mV (c) At what time is the emf zero?
Answer #1

i = 2t^{2} - 7t

The emf induced in an inductor is given by

\varepsilon = -L \frac{di}{dt}

\varepsilon = -0.07 \frac{d}{dt}( 2t^{2} - 7t)

\varepsilon = -0.07( 4t - 7)

a) t = 1 s

\varepsilon = -0.07( 4(1) - 7)

\varepsilon = 0.21 V

\varepsilon = 210 mV

b) t = 4s

\varepsilon = -0.07( 4(4) - 7)

\varepsilon = -0.63 V

The magnitude of emf is

\varepsilon = 0.63 V

\varepsilon =630 mV

c)

\varepsilon = -0.07( 4t - 7)

0 = -0.07( 4t - 7)

4t - 7= 0

t = \frac{7}{4}

t = 1.75 s